OpenStudy (anonymous):

Solve the following system of equations. Round your answers to the nearest tenth. Your answer will be in the form: (M,N). Find M + N, rounded to the nearest tenth.

OpenStudy (anonymous):

5x+2y=21 -2x+6y=-34